OpenStudy (anonymous):

please help, will fan and medal :) What is the value of the expression when p = –24 and q = 4? p/2q

OpenStudy (sissyedgar):

16 becuase we are goin to do the side with the number 1st. 2q ~ 2*4=8 {replace the variable with the number} p/8 ~ -24/8 = 16 {again, replace the variable with the number}

OpenStudy (anonymous):

@sissyedgar thank you so much :)

OpenStudy (sissyedgar):

no problem!

OpenStudy (anonymous):

@sissyedgar do you think you could check an answer for me?

OpenStudy (anonymous):

What is the value of g? -25/g=25 g = -625?

OpenStudy (anonymous):

never mind I got it, it was -1 :)